# Interpretable Graph Learning Over Sets of Temporally-Sparse Data

### v2

@@ Abstract
Real-world medical data often includes measurements from multiple signals that are collected at irregular and asynchronous time intervals. For example, different types of blood tests can be measured at different times and frequencies, resulting in fragmented and unevenly scattered temporal data. Similar issues of irregular sampling of different attributes occur in other domains, such as monitoring of large systems using event log files or the spread of fake news on social networks. Effectively learning from such data requires models that can handle sets of temporally sparse and heterogeneous signals. In this paper, we propose Graph Mixing Additive Networks (GMAN), a novel and interpretable-by-design model for learning over irregular sets of temporal signals. Our method achieves state-of-the-art performance in real-world medical tasks, including a $>\!4$ point increase in the AUROC score of in-hospital mortality prediction, compared to existing methods. We further showcase GMAN’s flexibility by applying it to a fake news detection task. We demonstrate how its interpretability capabilities, including node-level, graph-level, and subset-level importance, allow for transition phases detection, and gaining medical insights with real-world high-stakes implications. Finally, we provide theoretical insights on GMAN expressive power.
@@ 1 Introduction
Modern clinical data consist of diverse types of measurements, often collected at irregular and asynchronous time intervals. For example, a patient’s medical record often includes a set of blood tests taken over their lifetime, where each type of test is performed at its own frequency (e.g., inflammatory markers during one visit, liver function markers two years later, followed by a repeat of inflammatory markers that same year). As a result, the data can be effectively viewed as a set of sparse temporal trajectories. Similar patterns occur in other domains. For instance, the spread of a news article may follow distinct trajectories across platforms such as X (Twitter), Facebook, or WhatsApp, each governed by its own timing, user interactions, and propagation dynamics. Each such trajectory not only follows its own temporal dynamics but also reflects the underlying relational structure of the social network through which it propagates. As another example, software event logs often contain different types of events occurring at varying times and rates.
To learn prediction tasks over such data, conventional approaches often rely on imputation [7, 42, 47, 13] to flatten or regularize temporally irregular data into fixed-size representations. A common strategy is to align observations to a uniform time grid and fill in missing values using interpolation or learned imputation models. Alternatively, event-based data may be transformed into sequences of inter-event intervals or cumulative event counts. Such preprocessing imposes constraints that can potentially lose informative content, risks losing the conditional dependence structure of previous events, and ignore the natural set structure of the data, where each example contains a collection of related but non-synchronized temporal sequences.
In this paper, we introduce Graph Mixing Additive Networks (GMAN), a novel framework designed to learn directly from sets of sparse, irregular temporal trajectories. We handle irregular sampling by representing each trajectory as a directed graph and modelling the entire dataset as a set of such graphs, which may differ in size and the underlying information they represent. Our approach extends Graph Neural Additive Networks (GNAN) [4], which are interpretable by design but constrained by their restriction on non-linear feature interactions, limiting expressive power. In contrast, GMAN offers a flexible trade-off between interpretability and expressive power, overcoming GNAN’s constraints. Additionally, while GNAN operates only on single graphs, GMAN supports learning from sets of graphs, enabling further interpretability through analysis of the relative importance of individual graphs within the set.
We demonstrate the effectiveness of GMAN on real-world medical datasets, where it achieves state-of-the-art (SoTA) performance with substantial margins, while also providing valuable clinical and biological insights through its interpretability features. Additionally, we evaluate GMAN on a fake news detection task, showing that it generalizes well to sets of highly diverse and complex graphs and achieves SoTA results in this domain as well. Finally, we provide a theoretical analysis proving that GMAN is strictly more expressive than GNAN, and that the proposed mechanism for balancing interpretability and expressivity leads to strictly higher expressivity compared to models that enforce full interpretability.
@@
    We introduce GMAN, a novel interpretable and flexible method for learning over sets of sparse, irregular temporal trajectories.
@@
    We demonstrate the effectiveness of GMAN on real-world medical data and fake-news detection, providing SoTA performance, including a > 4 point increase in the AUROC score of predicting in-hospital mortality, and almost 2 points increase in the AUROC score of predicting Crohn’s Disease onset.
@@
    We demonstrate how GMAN can provide real clinical insights with high-stake implications, through various interpretability capabilities, including node-level, graph-level, and subset-level importance scores.
@@
    We provide a theoretical analysis on the expressive power of GMAN.
@@ 2 Related work
Graph Neural Networks (GNNs) [24, 16, 44, 50] have become the dominant framework for learning over graph-structured data, enabling flexible representation learning across diverse domains such as healthcare [31, 30, 33], chemistry [35, 23] and social networks [25, 39], amongst others. GNNs leverage both the graph topology and node features to compute learned representations for individual nodes or for entire graphs. Recently, Bechler-Speicher et al. [4] introduced Graph Neural Additive Networks (GNANs) a novel interpretable-by-design graph learning framework inspired by generalized additive models (GAMs) [19, 20]. GNAN applies univariate neural networks to each feature of the nodes separately, and then linearly combines their outputs across nodes to produce node-level and graph-level representations. As features are not mixed non-linearly, GNAN is fully interpretable, and provides feature-level and node-level interpretability which shows exactly how each feature and each node contribute to the final target variable.
Learning from irregularly sampled trajectories
@@
In many real-world settings, data is recorded as irregularly sampled trajectories with missing values from non-uniform timing and variable feature availability. Recurrent models [7] treat missing data as latent variables, while attention-based methods [13, 47] reconstruct them via contextual masking and temporal blocks. Diffusion models [42, 2, 38, 12] learn conditional distributions over missing values using score-based processes. Graph-based approaches use GNNs to model feature-time dependencies through bipartite graphs, adaptive message passing, or spatio-temporal attention [52, 11, 26, 51, 9]. Despite strong reconstruction, imputation pipelines often distort dynamics and may not improve prediction [34]. An alternative is to model sparsity directly. Neural and Latent ODEs [10, 36] address irregular gaps via continuous dynamics but are compute-intensive and rely on missingness encodings. Recent models [54] instead use graph representations to capture temporal sparsity without imputation.
3 Graph Mixing Additive Networks
Figure 1: Overview of GMAN architecture. Each input is a set of graphs, an ExtGNAN is applied to each graph to obtain graph-level representations. Graphs can optionally be grouped and aggregated using a DeepSet module, allowing non-linear combination of related signals. This design enables a flexible trade-off between interpretability and expressivity.
In this section, we present GMAN, an interpretable and flexible method for effectively learning over sets of trajectories of varying size. Let $S=\{G_{1},\dots,G_{m}\}$ be a set of $m$ graphs. Each node $v$ is associated with a feature vector $x_{v}\in\mathbb{R}^{d}$ and a time-stamp $t_{v}$. For instance, in the case of a patient’s blood test data, each graph corresponds to a specific biomarker, and each node within the graph represents an individual measurement of that biomarker, annotated with its feature vector and time of collection. We denote the entry $c$ of a vector $\mathbf{h}$ by $[\mathbf{h}]_{c}$, and the set of entries corresponding to a set of features $S$ by $[\mathbf{h}]_{S}$.
We assume that the graphs in $S$ are partitioned into $k,1\leq k\leq m$ disjoint subsets $S_{k}$ such that $\bigcup_{i=1}^{k}S_{i}=S$. The partition $\{S_{i}\}_{i=1}^{k}$ provides a flexible trade-off between expressivity and interpretability. GMAN linearly aggregates representations of the subsets of $S$ to form a final set representation, and then assigns a single label to $S$. The level of interpretability that GMAN provides for each graph depends on the size of the subset it belongs to. When a subset contains a single graph, GMAN offers fine-grained, node-level importance scores. In contrast, for larger subsets, it provides only set-level importance scores—trading interpretability for improved expressivity. This design enables a flexible trade-off between interpretability and expressive power, controlled by the chosen partitioning strategy.
@@ 3 Graph Mixing Additive Networks
A visual overview of GMAN is presented in Figure 1. First, GMAN applies a function $\Phi_{i}$ to each subset $S_{i}$ to obtain a representation of the subset $S_{i}$, denoted as $\mathbf{h}_{i}\in\mathbb{R}^{d}$.
@@
Then, it produce a representation for the whole set, $\mathbf{h}_{S}$ by summing the subsets’ $\mathbf{h}_{S}=\sum_{i=1}^{k}h_{i}$. Finally, to produce the label, it sums over the $d$ entries of $\mathbf{h}_{S}$. Overall:
$\textsc{GMAN}(S)=\sum_{c=1}^{d}\sum_{i=1}^{k}[\Phi_{i}(S_{i})]_{c}$ (1)
@@
For subsets of size one, $\Phi_{i}(S_{i})$ applies an Extended GNAN (ExtGNAN), as described in Section 3.1. For subsets containing multiple graphs, a featuregroupgnan is applied to each graph, followed by a DeepSet aggregation [53] over the resulting vectors. Importantly, each subset is assigned its own ExtGNAN, and all graphs within a subset share the same one. A DeepSet first applies a neural network $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for each vector in the set $\{h_{l}\}_{G_{l}\in S_{i}}$, sums the results, and then applies another neural network $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$.
@@
Here, $g$ and $f$ are neural networks of arbitrary depth and width. We now turn to define ExtGNAN.
@@ 3.1 ExtGNAN
In GNAN, univariate neural networks are applied to each feature of each node in isolation, to learn a representation for a graph. This has the benefit of generating interpretable models as features do not mix non-linearly. Nonetheless, when interactions between features are crucial for the task, it may result in sub-par performance. Therefore, ExtGNAN extends GNAN by allowing multivariate neural networks to operate on groups of features to gain accuracy at the cost of reducing the feature-level interpretability of the model.
Assume that the features are partitioned into $K$ subsets $\{F_{l}\}_{l=1}^{K}$. For any subset of features greater than one, ExtGNAN applies a multivariate neural network for all the features in the subset together, instead of a univariate neural network for each one separately. To learn a representation of a graph $G$, ExtGNAN first computes representations for the nodes of $G$ as follows.
ExtGNAN learns a distance function $\rho(x;\theta):\mathbb{R}\rightarrow\mathbb{R}$ and a set of feature shape functions $\{\psi_{l}\}_{l=1}^{K},\psi_{l}(X;\theta_{k}):\mathbb{R}^{|F_{l}|}\rightarrow\mathbb{R}^{|F_{l}|}$. Each of these functions is a neural network of arbitrary depth. For brevity, we omit the parameterization $\theta$ and $\theta_{k}$ for the remainder of this section.
The entries of the representation of node $j$ corresponding to the indices of the features in $F_{l}$, denoted as $[\mathbf{h}_{j}]_{F_{l}}$, is computed by summing the contributions of the features in the subset $F_{l}$ from all nodes in the graph:
$[\mathbf{h}_{j}]_{F_{l}}=\sum_{w\in V}\rho\left(\Delta(w,j)\right)\cdot\psi_{l}\left([\mathbf{X}_{w}]_{F_{l}}\right),$
where $\Delta(w,j)=t_{w}-t_{j}$ and $[\mathbf{X}_{w}]_{F_{l}}$ are the features of node $w$ corresponding to the subset $F_{l}$.
@@
This concludes the description of ExtGNAN, which computes graph-level representations. Next, we describe how GMAN combines these representations across sets and enables multi-level interpretability.
3.2 Node, graph and set importance
GMAN retains all interpretability properties of GNAN, including feature-level and node-level importance. However, it extends beyond GNAN by operating on sets of graphs rather than single graphs, enabling additional forms of interpretability such as graph-level and subset-level importance. Because GMAN allows a flexible trade-off between interpretability and expressivity, permitting non-linear mixing within graph subsets, some adaptations are required to extract meaningful attributions under this more expressive regime. Due to space constraints, we present only the formulations for node and graph importance here; additional interpretability measures are provided in the Appendix. Section 4 illustrates how node-level attributions yield meaningful real-world insights.
We can extract the total contribution of each node $j$ to the prediction by summing the contributions of the node across all feature sets. This is only valid when the node belongs to a graph that is not combined non-linearly with other graphs, i.e., it belongs to a subset of size one.

### v3

# SuperMAN: Interpretable and Expressive Networks over Temporally Sparse Heterogeneous Data

@@ Abstract
Real-world temporal data often consists of multiple signal types recorded at irregular, asynchronous intervals. For instance, in the medical domain, different types of blood tests can be measured at different times and frequencies, resulting in fragmented and unevenly scattered temporal data. Similar issues of irregular sampling occur in other domains, such as the monitoring of large systems using event log files. Effectively learning from such data requires handling sets of temporally sparse and heterogeneous signals. In this work, we propose Super Mixing Additive Networks (SuperMAN), a novel and interpretable-by-design framework for learning directly from such heterogeneous signals, by modeling them as sets of implicit graphs. SuperMAN provides diverse interpretability capabilities, including node-level, graph-level, and subset-level importance, and enables practitioners to trade finer-grained interpretability for greater expressivity when domain priors are available. SuperMAN achieves state-of-the-art performance in real-world high-stakes tasks, including predicting Crohn’s disease onset and hospital length of stay from routine blood test measurements and detecting fake news. Furthermore, we demonstrate how SuperMAN’s interpretability properties assist in revealing disease development phase transitions and provide crucial insights in the healthcare domain.
@@ 1 Introduction
Modern clinical data consist of diverse types of signals, often collected at irregular and asynchronous time intervals. For example, a patient’s medical record may include a set of blood tests taken over their lifetime, where each type of test is performed at its own frequency. As a result, the data can be viewed as a set of sparse temporal signals, each with its own fragmented temporal structure. Similar patterns occur in other domains. For instance, the spread of news articles in social media networks often follow asynchronous tree-like patterns of dissemination. Another example is system event logs, which typically include different types of events occurring at varying times and rates.
A common approach for learning from such data is to align the signals to a fixed-size time grid, thereby enforcing a shared timeline. This is typically achieved by trimming or aggregating signals and filling in missing values through interpolation or learned imputation models (Cao et al., 2018; Tashiro et al., 2021; Wu et al., 2022; Du et al., 2023). However, these procedures can lead to substantial information loss and ignore the informative patterns in the irregularity itself, such as the varying time intervals between different measurement types.
In this work, we introduce Super Mixing Additive Networks (SuperMAN), a novel framework designed to learn directly from heterogeneous, temporally sparse and irregular signals, without information loss or imputation. SuperMAN models these signals as sets of implicit graphs by introducing a distance function between each pair of signals within each type of signal. For example, in blood test data, each biomarker type can be modeled as a directed graph whose nodes correspond to individual measurements, with the distance between each pair of measurements being the time delta between them. These implicit graphs may differ in structure, size, and feature space, reflecting the diversity of real-world heterogeneous signals.
SuperMAN allows for multi-resolution interpretability capabilities, including node-level, graph-level, and subset-level importance. It also enables practitioners to trade fine-grained interpretability for greater expressivity when domain priors are available. Specifically, signal graphs can be grouped into subsets, allowing for more expressive modeling of non-linear interactions within the subsets. This shifts interpretability from individual nodes or graphs to the subset level. SuperMAN builds on Graph Neural Additive Networks (GNAN) (Bechler-Speicher et al., 2024), an interpretable class of GNNs that operate on individual nodes or graphs. In contrast, SuperMAN is expressly designed for sets of implicit graphs and introduces a graph-grouping mechanism that aggregates signals within subsets of graphs while preserving an additive decomposition across subsets. Rather than applying GNAN directly, SuperMAN employs an extended and more flexible variant, which we denote by ExtGNAN. Within each graph, ExtGNAN generalises GNAN’s univariate feature-shape networks to multivariate feature groups, allowing non-linear dependencies among related features while retaining additive transparency at the group level. In analogy to signal grouping, ExtGNAN supports grouping features into subsets, thereby replacing feature-level interpretability with subset-level interpretability, such as importance scores. We prove that grouping signals or features strictly increases expressivity, and that SuperMAN is strictly more expressive than GNAN.
We demonstrate the effectiveness of SuperMAN on real-world high-stakes medical datasets, where it achieves state-of-the-art (SoTA) performance while also providing valuable clinical and biological insights through its interpretability. In addition, we show that SuperMAN attains SoTA performance in fake news detection, highlighting its flexibility in operating on sets of graphs with arbitrary structure. This contrasts with existing approaches, which are restricted to sets of path-like signals and offer no interpretability. Finally, we demonstrate how SuperMAN’s interpretability reveals phase transitions and provides crucial insights in healthcare.
@@
    We introduce SuperMAN, a novel framework for learning directly from sets of sparse, irregular temporal signals, without information loss or imputation.
@@
    We allow practitioners to integrate domain priors, when available, by grouping features or signal types into subsets. This shifts interpretability from fine-grained (node- or feature-level) to subset-level, while strictly improving expressivity. This capability is particularly valuable in the medical domain, where such priors are common.
@@
    We provide a theoretical analysis proving that groupings of features and signals make SuperMAN strictly more expressive.
@@
    We demonstrate the effectiveness of SuperMAN on real-world high-stakes medical datasets and fake news detection, achieving state-of-the-art performance in both domains.
  5. 5.
    We show that SuperMAN provides valuable interpretability capabilities, including node-level, graph-level, and subset-level importance, which yield meaningful clinical and biological insights.
@@ 2 Related work
Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Gilmer et al., 2017; Velickovic et al., 2017; Xu et al., 2018) have become the dominant framework for learning over graph-structured data, enabling flexible representation learning across diverse domains such as healthcare (Paul et al., 2024; Ochoa and Mustafa, 2022; Peng et al., 2023), chemistry (Reiser et al., 2022; Jumper et al., 2021) and social networks (Li et al., 2023; Sharma et al., 2024), among others. GNNs leverage both the graph topology and node features to compute learned representations for individual nodes or for entire graphs. Recently, Bechler-Speicher et al. (2024) introduced Graph Neural Additive Networks (GNANs), a novel interpretable-by-design graph learning framework inspired by generalized additive models (GAMs) (Hastie and Tibshirani, 1986; 1987). GNAN applies univariate neural networks to each feature of the nodes separately, and then linearly combines their outputs across nodes to produce node-level and graph-level representations. As features are not mixed non-linearly, GNAN is fully interpretable, and provides feature-level and node-level interpretability which shows exactly how each feature and each node contributes to the final target variable.
Learning from sparse data
@@
In many real-world settings, data often presents missing values from irregular sampling and variable feature availability. Recurrent models (Cao et al., 2018) treat missing data as latent variables, while attention-based methods (Du et al., 2023; Wu et al., 2022; Tipirneni and Reddy, 2022; Labach et al., 2023) reconstruct them via contextual masking and temporal blocks. Diffusion models (Tashiro et al., 2021; Alcaraz and Strodthoff, 2022; Senane et al., 2024; Dai et al., 2024) learn conditional distributions over missing values using score-based processes. Graph-based approaches use GNNs to model feature dependencies through bipartite graphs, adaptive message passing, or spatio-temporal attention (You et al., 2020; Cini et al., 2022; Marisca et al., 2022; Ye et al., 2021; Chen et al., 2024). These imputation methods often distort dynamics and may not improve prediction (Qian et al., 2025). An alternative is to model sparsity directly. Neural and Latent ODEs (Chen et al., 2018; Rubanova et al., 2019) address irregular gaps via continuous dynamics but are compute-intensive and rely on missingness encodings. Recent models (Zhang et al., 2021) have used graph representations to capture sparsity without imputation. Importantly, none of the aforementioned methods offer built-in multi-grain interpretability in the way that SuperMAN does.
3 Super Mixing Additive Networks
Figure 1: In this example, the input is a set of three graphs, $G_{1},G_{2},G_{3}$, grouped into two subsets $S_{1}$ and $S_{2}$. Within each subset, the same ExtGNAN instance is applied to all graphs to produce their graph-level representations. For subsets containing multiple graphs, a DeepSets module aggregates these graph representations into a single subset representation. For subsets of size one, the subset representation is simply the graph representation itself. The final set representation is then obtained by summing the subset representations, and the final label prediction is produced by summing the entries of this set representation.
In this section, we present SuperMAN, an interpretable and flexible method for learning over sets of arbitrary graphs.
Preliminaries
SuperMAN acts on a set of $m$ graphs $S=\{G_{1},\dots,G_{m}\}$, which can be directed or undirected. Each node $v\in G_{i}$, $1\leq i\leq m$ is associated with a feature vector $x_{v}\in\mathbb{R}^{d}$ and a time-stamp $t_{v}$. SuperMAN utilizes the distances between each pair of nodes in each graph, denoted by $\Delta_{uv}$, where
[…]
Graphs may be provided directly through an explicit layout, e.g., the news propagation graphs we evaluate SuperMAN over in  section˜4. More commonly, however, sparse temporal heterogeneous data does not come with a predefined graph structure. In such cases, we construct directed path-graphs for each signal type based on its measurement time stamps, as demonstrated for the medical dataset in section˜4. For instance, in patient blood test records, each graph corresponds to a specific biomarker. Nodes in each graph represent individual measurements of that biomarker, annotated with the observed test value as a feature and the associated time stamp. We mark vectors with bold, and denote the entry $c$ of a vector $\mathbf{h}$ by $\mathbf{h}]_{c}$, and the set of entries corresponding to a set of features $S$ by $\mathbf{h}]_{S}$.
Signal Grouping
To incorporate domain priors, the graphs in $S$ can be partitioned into $k$ disjoint subsets $S_{1},\dots,S_{k}$ with $\bigcup_{i=1}^{k}S_{i}=S$. If at least one subset $S^{\prime}$ contains multiple graphs, the model’s expressivity increases, as we prove in Theorem˜3.2, at the cost of shifting interpretability from individual nodes and graphs in $S^{\prime}$ to the subset level. In practice, this means we can attribute importance to $S^{\prime}$ as a whole, but not to its individual components. This trade-off—enhanced expressivity at reduced granularity of interpretability is especially valuable in domains such as medicine, where priors often suggest natural groupings of signals and interpretability is only needed at the subset level. This is demonstrated in Section˜4.
SuperMAN linearly aggregates representations of the subsets of $S$ to form a final set representation, and then assigns a single label to $S$.
@@ 3 Super Mixing Additive Networks
First, SuperMAN applies a function $\Phi_{i}$ to each subset $S_{i}$ to obtain a representation of the subset $S_{i}$, denoted as $\mathbf{h}_{i}\in\mathbb{R}^{d}$.
@@
Then, it produces a representation for the whole set, $\mathbf{h}_{S}$ by summing the subsets’ $\mathbf{h}_{S}=\sum_{i=1}^{k}h_{i}$. Finally, to produce the label, it sums over the $d$ entries of $\mathbf{h}_{S}$. Overall:
$\textsc{SuperMAN}(S)=\sum_{c=1}^{d}\sum_{i=1}^{k}[\Phi_{i}(S_{i})]_{c}$ (1)
@@
For subsets of size one, $\Phi_{i}(S_{i})$ applies an Extended GNAN (ExtGNAN), as described in Section˜3.1. For subsets containing multiple graphs, a featuregroupgnan is applied to each graph, followed by a DeepSet aggregation (Zaheer et al., 2018) over the resulting vectors. Importantly, each subset is assigned its own ExtGNAN, and all graphs within a subset share the same one. A DeepSet first applies a neural network (NN) $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for each vector in the set $\{h_{l}\}_{G_{l}\in S_{i}}$, sums the results, and then applies another NN $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$.
@@
Here, $g$ and $f$ are NNs of arbitrary depth and width. A high-level visual overview of SuperMAN is presented in Figure˜1. We now turn to define ExtGNAN.
@@ 3.1 ExtGNAN
In GNAN, univariate NNs are applied to each feature of each node in isolation, to learn a representation for a graph. This has the benefit of generating interpretable models as features do not mix non-linearly. Nonetheless, when interactions between features are crucial for the task, or feature-level interpretability is not required for all features, it may result in sub-par performance. Therefore, ExtGNAN extends GNAN by allowing multivariate NNs to operate on groups of features to gain accuracy at the cost of reducing the feature-level interpretability only for features that are grouped together, and obtaining interpretability for their subset as a whole instead.
Assume that the features are partitioned into $K$ subsets $\{F_{l}\}_{l=1}^{K}$. For any subset of features greater than one, ExtGNAN applies a multivariate NN for all the features in the subset together, instead of a univariate NN for each one separately. To learn a representation of a graph $G$, ExtGNAN first computes representations for the nodes of $G$ as follows.
ExtGNAN learns a distance function $\rho(x;\theta):\mathbb{R}\rightarrow\mathbb{R}$ and a set of feature shape functions $\{\psi_{l}\}_{l=1}^{K},\psi_{l}(X;\theta_{k}):\mathbb{R}^{|F_{l}|}\rightarrow\mathbb{R}^{|F_{l}|}$. Each of these functions is a NN of arbitrary depth. For brevity, we omit the parameterization $\theta$ and $\theta_{k}$ for the remainder of this section.
The entries of the representation of node $j$ corresponding to the indices of the features in $F_{l}$, denoted as $\mathbf{h}_{j}]_{F_{l}}$, is computed by summing the contributions of the features in the subset $F_{l}$ from all nodes in the graph:
$\mathbf{h}_{j}]_{F_{l}}=\sum_{w\in V}\rho\left(\Delta(w,j)\right)\cdot\psi_{l}\left([\mathbf{X}_{w}]_{F_{l}}\right),$
where $\Delta(w,j)=t_{w}-t_{j}$ and $\mathbf{X}_{w}]_{F_{l}}$ are the features of node $w$ corresponding to the subset $F_{l}$.
@@
This concludes the description of ExtGNAN, which computes graph-level representations. We provide a complexity analysis of SuperMAN in the appendix, where we also discuss how $\rho$ can be masked to adapt to any complexity limitation, including a linear one. Next, we describe how SuperMAN combines these representations across sets and enables multi-level interpretability.
3.2 Node, graph and subset importance
SuperMAN retains all interpretability properties of GNAN, including feature-level and node-level importance. However, it extends beyond GNAN by operating on sets of graphs rather than single graphs, enabling additional forms of interpretability such as graph-level and subset-level importance. Because SuperMAN allows a flexible trade-off between interpretability and expressivity, permitting non-linear mixing within graph subsets, some adaptations are required to obtain importance scores. A key property of SuperMAN’s interpretability is that its importance scores directly reflect the contribution of each node, graph, or subset to the predicted label, since these terms are combined additively to produce the final output. Section˜4 illustrates how node-level importances yield insightful real-world insights.
We can extract the total contribution of each node $j$ to the prediction by summing the contributions of the node across all feature subsets. This is only valid when the node belongs to a graph that is not combined non-linearly with other graphs, i.e., it belongs to a subset of size one.